# Emergence and Evolution of Interpretable Concepts in Diffusion Models

### v1

@@ Abstract
Diffusion models have become the go-to method for text-to-image generation, producing high-quality images from noise through a process called reverse diffusion. Understanding the dynamics of the reverse diffusion process is crucial in steering the generation and achieving high sample quality. However, the inner workings of diffusion models is still largely a mystery due to their black-box nature and complex, multi-step generation process. Mechanistic Interpretability (MI) techniques, such as Sparse Autoencoders (SAEs), aim at uncovering the operating principles of models through granular analysis of their internal representations. These MI techniques have been successful in understanding and steering the behavior of large language models at scale. However, the great potential of SAEs has not yet been applied toward gaining insight into the intricate generative process of diffusion models. In this work, we leverage the SAE framework to probe the inner workings of a popular text-to-image diffusion model, and uncover a variety of human-interpretable concepts in its activations. Interestingly, we find that even before the first reverse diffusion step is completed, the final composition of the scene can be predicted surprisingly well by looking at the spatial distribution of activated concepts. Moreover, going beyond correlational analysis, we show that the discovered concepts have a causal effect on the model output and can be leveraged to steer the generative process. We design intervention techniques aimed at manipulating image composition and style, and demonstrate that (1) in early stages of diffusion image composition can be effectively controlled, (2) in the middle stages of diffusion image composition is finalized, however stylistic interventions are effective, and (3) in the final stages of diffusion only minor textural details are subject to change.
@@ 1 Introduction
Figure 1: Coarse image composition emerges during the very first generation step in the diffusion process. We generate an image with the prompt "Several men walking on the dirt with palm trees in the background". Our interpretability framework can predict segmentation masks for each object mentioned in the input prompt, solely relying on model activations cached during the first diffusion step. At this early stage, the posterior mean predicted by the diffusion model does not contain any visual clues about the final generated image.
@@
  • What level of image representation is present in the early, ’chaotic’ stage of the generative process?
@@ 2 Background
Interpreting diffusion models – There has been significant effort towards interpreting diffusion models. Authors in Tang et al. [44] find that the cross-attention layers in diffusion models with a U-Net backbone – such as SDXL [32] and Stable Diffusion [34] – can be used to generate saliency maps corresponding to textual concepts. Another line of work focuses on finding interpretable editing directions in diffusion U-Nets to control the image generation process. In particular, Kwon et al. [23] and Haas et al. [13] focus on manipulating bottleneck features, Park et al. [31] finds edit directions based on the SVD of the Jacobian between the input and bottleneck layer of the U-Net, Chen et al. [5] considers the Jacobian between the input and the posterior mean estimate rather than the bottleneck, Orgad et al. [30], Gandikota et al. [10] modify the key and value projection matrices, and Epstein et al. [9], Chen et al. [4] seek to control object position, size, shape directly by thresholding attention maps.
In recent work [43], authors train SAEs on the activations of a distilled, single-step diffusion model [37] (SDXL Turbo). In particular, they target certain cross-attention transformer blocks in the U-Net and train SAEs based on the residual update made by the transformer block. The features in the latent space of SAEs are found to be highly interpretable. Our work differs from theirs in two important ways. First, we analyze the time-evolution of interpretable concepts during the generative process, a key component in understanding and controlling the diffusion process, which is not captured by a single-step model. Second, they leverage vision-language foundation models to extract the semantics of SAE features by reasoning about and summarizing the commonalities between groups of images that activate specific features. This technique, however, is difficult to scale to large number of images due to the limited context of such models and is afflicted by the reasoning limitations, biases and hallucinations of the foundation model. In contrast, we propose a simple, scalable pipeline to extract interpretations for SAE features in the form of a flexible concept dictionary, leveraging open-set object detectors and segmentation models. Concurrent work [7] demonstrates the potential of SAEs in machine unlearning for diffusion models. Even though, similar to our work, they study a non-distilled diffusion model, their analysis focuses on identifying and removing particular concepts from generated images, and not on understanding the time-evolution of internal representations. In fact, they train a single SAE jointly for all time steps, whereas we perform a more granular analysis and train separate SAEs specialized to each time step. Recent work [20] leverages the SAE framework for controlled text-to-image generation. Different from our work, they train SAEs on the activations of the separate text-encoder that guides the diffusion model, and thus they do not investigate the visual representations of the diffusion model itself.
@@ 3.1 SAE Architecture and Loss
In this section, we discuss the design choices behind our SAE model. We opt for $k$-sparse autoencoders (with TopK activation) given their success with GPT-4 [11] and SDXL Turbo [43]. In particular, let $\bm{x}\in\mathbb{R}^{d}$ denote the input activation to the autoencoder that we want to decompose into a sparse combination of features. Then, we obtain the latent $\bm{z}\in\mathbb{R}^{n_{f}}$ by encoding $\bm{x}$ as
$\bm{z}=\mathcal{E}\left(\bm{x}\right)=\text{TopK}\left(\text{ReLU}\left(\bm{W}_{enc}\left(\bm{x}-\bm{b}\right)\right)\right),$
where $\bm{W}_{enc}\in\mathbb{R}^{n_{f}\times d}$ denotes the learnable weights of the encoder, $\bm{b}\in\mathbb{R}^{d}$ is a learnable bias term, and TopK function keeps the top $k$ highest activations and sets the remaining ones to 00. Note, that due to the superposition hypothesis, we wish the encoding to be expansive and therefore $n_{f}\gg d$. Then, a decoder is trained to reconstruct the input from the latent $\bm{z}$ in the form
$\hat{\bm{x}}=\mathcal{D}\left(\bm{z}\right)=\bm{W}_{dec}\bm{z}+\bm{b},$
where $\bm{W}_{dec}\in\mathbb{R}^{d\times n_{f}}$ represents the learnable weights of the decoder. Note, that the bias term is shared between the encoder and decoder. We refer to $\bm{f}_{i}=\bm{W}_{dec}[:,i]$ columns of $\bm{W}_{dec}$ as concept vectors. We obtain the learnable parameters by optimizing the reconstruction error
where $\alpha$ is a hyperparameter.
@@ 3.2 Collecting Model Activations
In this work, we use Stable Diffusion v1.4 (SDv1.4) [34] as our diffusion model due to its widespread use. Inspired by Surkov et al. [43], we use 1.5M training prompts from the LAION-COCO dataset [38] and store $\Delta_{\ell,t}\in\mathbb{R}^{H_{\ell}\times W_{\ell}\times d_{\ell}}$, the difference between the output and input of the $\ell$th cross-attention transformer block at diffusion time $t$ (i.e. the update to the residual stream). We train our SAE to reconstruct features individually along the spatial dimension. That is the input to the SAE is $\Delta_{\ell,t}[i,j,:]$ for different spatial locations $(i,j)$ whereas $\ell$ and $t$ are fixed and to be specified next.
To capture the time-evolution of concepts, we collect activations at timesteps corresponding to $t\in\left[0.0,0.5,1.0\right]$ and analyze final ($t=0.0$, close to final generated image), middle, and early ($t=1.0$, close to pure noise) diffusion dynamics respectively. For each timestep $t$, we target 3 different cross-attention blocks in the denoising model of SDv1.4: down_blocks.2.attentions.1, mid_block.attentions.0, up_blocks.1.attentions.0. We refer to these as down_block, mid_block, up_block for brevity. We specifically include the mid_block or the bottleneck layer of the U-Net since earlier work found interpretable editing directions here [23]. Other blocks are chosen to be the closest to the bottleneck layer in the downsampling and upsampling paths of the U-Net. The performance of text guidance is improved through Classifier-Free Guidance (CFG) [14]. The model output is modified as $\tilde{\varepsilon}_{\bm{\theta}}(\bm{x}_{t},t,\bm{c})=\varepsilon_{\bm{\theta}}(\bm{x}_{t},t,\bm{c})+\omega\left(\varepsilon_{\bm{\theta}}(\bm{x}_{t},t,\bm{c})-\varepsilon_{\bm{\theta}}(\bm{x}_{t},t,\bm{\varnothing})\right),$ where $\omega$ denotes the guidance scale, $\bm{c}$ is the conditioning input and $\bm{\varnothing}$ is the null-text prompt. At each timestep we collect both the text-conditioned diffusion features (called cond) and null-text-conditioned features (denoted by uncond).
To provide a granular and in-depth analysis, we train separate SAEs for different block, conditioning and timestep combinations. Training results are in Appendix A. In this work, we focus on cond features, as we hypothesize that they may be more aligned with human-interpretable concepts due to the direct influence of language guidance through cross-attention (more on this in Appendix C).
@@ 3.3 Extracting interpretations from SAE features
In particular, we represent each concept by an associated list of objects, constituting a concept dictionary. The keys are unique concept identifiers (CIDs) assigned to each of the concept vectors of the SAE. The values correspond to objects that commonly occur in areas where the concept is activated. To build the concept dictionary (Figure 3), we first sample a set of text prompts, generate the corresponding images using a diffusion model and extract the SAE activations for each CID during generation. We obtain ground truth annotations for each generated image using a pre-trained vision pipeline, that combines image tagging, object detection and semantic segmentation, resulting in a mask and label for each object in generated images. Finally, we evaluate the alignment between our ground truth masks and the SAE activations for each CID, and assign the corresponding label to the CID only if there is sufficient overlap.
@@
Figure 2: Curating the concept dictionary: 1) We cache SAE activations for various time steps and blocks during image generation. 2) We leverage a pipeline of image tagging, open-set object detection and promptable segmentation to annotate the generated image with segmentation masks and corresponding object labels. 3) We find SAE activations that sufficiently overlap with the object masks. 4) We add the overlapping object’s label to the concept dictionary under the matching SAE activation’s CID.
Figure 3: Predicting image composition: 1) We cache SAE activations during the very first diffusion step (or other time step of interest) and extract top activated concepts for each spatial location. 2) For each spatial location, we fetch the associated objects from the concept dictionary and produce a conceptual embedding via Word2Vec. 3) We compare the conceptual embedding at each location to the target word embeddings from the input prompt and predict a segmentation map based on cosine similarity.
@@ 3.4 Predicting image composition from SAE features
Leveraging the concept dictionary, we predict the final image composition based on SAE features at any time step (Figure 3), allowing us to gain invaluable insight into the evolution of image representations in diffusion models. Suppose that we would like to predict the location of a particular object in the final generated image, but before the reverse diffusion process is completed. First, given SAE features from a given intermediate time step, we extract the top activating concepts for each spatial location. Next, we create a conceptual map of the image by assigning a word embedding to each spatial location based on our curated concept dictionary. This conceptual map shows how image semantics, described by localized word embeddings, vary spatially across the image. Given a concept we would like to localize, such as an object from the input prompt, we produce a target word embedding and compare its similarity to each spatial location in the conceptual map. To produce a predicted segmentation map, we assign the target concept to spatial locations with high similarity, based on a pre-defined threshold value. This technique can be applied to each object present in the input prompt (or to any concepts of interest) to predict the composition of the final generated image.
@@ 3.5 Causal intervention techniques
Figure 4: An overview of our SAE intervention technique. The prompt "An apple in a basket" specifies the necessary concepts but is vague in terms of spatial composition. We intercept activations of the denoising model and edit the latents after encoding them with the SAE. For the features that are spatially located in the bottom-right quadrant, we increase the coefficient corresponding to "apple" concept. For the remaining features, latents corresponding to "apple" concept are set to 00. After the intervention, generated image satisfies the specified layout where all the apples are located in the bottom-right quadrant.
Spatially targeted interventions – To assess layout controllability using the discovered concepts, we propose a simple task: enforce a specific object to appear only in a designated quadrant (e.g., top-left) of the image. To achieve this, we intercept activations and edit features in the SAE latent space by amplifying the desired concept in the target region and setting it to 00 otherwise. Recall, that the contribution of the $\ell$th transformer block at time $t$ is given by $\Delta_{\ell,t}$. Let $\bm{Z}_{\ell,t}$ denote the latents after encoding the activations with the SAE encoder $\mathcal{E}$. Let $S$ denote the set of coordinates to which we would like to restrict the object. Let $C_{o}$ be the set of CIDs that are relevant to object $o$. We wish to modifty the latents as follows:
@@
where $\beta$ is our intervention strength. However, decoding the modified latents directly is suboptimal as the SAE cannot reconstruct the input perfectly. Instead, we modify the activations directly using the concept vectors. The modification in Eq. (1) can be equivalently written as:
@@
An overview of this intervention can be seen in Eq. (4). In prior experiments, we observe that the same intervention strength $\beta$ does not work well across different objects $o$. To solve this, we introduce a normalization where the intervention at a spatial coordinate $\left(i,j\right)$ is proportional to the norm of the latent at that coordinate $\left\|\bm{Z}_{\ell,t}[i,j]\right\|$. Therefore, the effective intervention strength is $\beta_{ij}=\beta\left\|\bm{Z}_{\ell,t}[i,j]\right\|$.
Global interventions – Beyond image composition, we investigate whether image style can be manipulated through our discovered concepts. To this end, given a CID $c$ related to the style of interest, as image style is a global property we modify the activation at each spatial location as follows:
@@ 4 Experiments
(a) $t=1.0$
(b) $t=0.0$
Figure 5: Concept dictionary and visualization of the activation maps for the top $5$ activating concepts extracted from up_blocks.1.attentions.0 for the (a) first and (b) last diffusion steps. Sample ID: $2000031$
@@ 4.1 Building the concept dictionary
We sample $40k$ prompts from the LAION-COCO dataset from a split that has not been used to train the SAEs. We build the concept dictionary following our technique introduced in Section 3.3. For annotating generated images, we leverage RAM [47] for image tagging, Grounding DINO [25] for open-set object detection and SAM [21] for segmentation, following the pipeline in Ren et al. [33]. We assign a label to a specific CID if the IoU between the corresponding annotated mask and activation is greater than $0.5$. We binarize the activation map for the IoU calculation by first normalizing to $[0,1]$ range, then thresholding at $0.1$. We visualize the top $5$ activating concepts and the corresponding concept dictionary entries in Figure 5(b) for a generated sample. More samples can be found in Appendix E.
4.2 Qualitative analysis of concept activations
We visualize the activation maps for top $10$ (in terms of mean activation across the spatial dimensions) activating concepts in Figure 6 across time steps. Based on our empirical observations, the activations can be grouped into the following categories.
Local semantics – Most concepts fire in semantically homogeneous regions, producing a semantic segmentation mask for a particular concept. Examples include the segmentation of the plate, food items and background in Figure 6. We observe that these semantic concepts can be redundant in the sense that multiple concepts often fire in the same region (e.g. see Fig. 6, second row with multiple concepts focused on the food in the bowl). We hypothesize that these duplicates may add different conceptual layers to the same region (e.g. food and round in the previous example). In terms of diffusion time, we observe that the segmentation masks are increasingly more accurate with respect to the final generated image, which is expected as the final scene progressively stabilizes during the diffusion process. This observation is more thoroughly verified in Section 4.3 and Figure 7(a). In terms of different U-Net blocks, we observe that up_block provides the most accurate segmentation of the final scene, especially at earlier time steps.
Context-free – We observe that some concepts fire exclusively in specific, structured regions of the image, such as particular corners or bordering edges of the image, irrespective of semantics. We hypothesize that these concepts may be a result of optimization artifacts, and are leveraged as semantic-independent knobs for the SAE to reduce reconstruction error. Visual examples and further discussion can be found in Appendix F.
More visualized concept activations for multiple blocks, time steps and samples can be found in Appendix D.
@@ 4.4 Effectiveness of interventions across diffusion time
Spatially targeted interventions– We consider bee, book, and dog as the objects of interest and attempt to restrict them to four different quadrants: top-left, top-right, bottom-left, and bottom-right. In order to find the CIDs to be intervened on, we sweep the concept dictionary of the given time step and collect all the CIDs where the word of interest appears. Results are summarized in Figure 8.
Global interventions– Through our concept dictionary and visual inspection of top dataset examples at $t=0.5$, we select the following CIDs: #$1722$ that controls the cartoon look of the image, #$524$ appears mostly with beach images where sea and sand are visible together, and #$2137$ activates the most on paintings (top activating images can be found in Appendix G). We find matching concepts for other time steps by picking the CIDs with the highest Word2Vec embedding similarity to the above target CIDs. An overview of results is depicted in Figure 9.

### v2

@@ Abstract
Diffusion models have become the go-to method for text-to-image generation, producing high-quality images from pure noise. However, the inner workings of diffusion models is still largely a mystery due to their black-box nature and complex, multi-step generation process. Mechanistic interpretability techniques, such as Sparse Autoencoders (SAEs), have been successful in understanding and steering the behavior of large language models at scale. However, the great potential of SAEs has not yet been applied toward gaining insight into the intricate generative process of diffusion models. In this work, we leverage the SAE framework to probe the inner workings of a popular text-to-image diffusion model, and uncover a variety of human-interpretable concepts in its activations. Interestingly, we find that even before the first reverse diffusion step is completed, the final composition of the scene can be predicted surprisingly well by looking at the spatial distribution of activated concepts. Moreover, going beyond correlational analysis, we design intervention techniques aimed at manipulating image composition and style, and demonstrate that (1) in early stages of diffusion image composition can be effectively controlled, (2) in the middle stages image composition is finalized, however stylistic interventions are effective, and (3) in the final stages only minor textural details are subject to change. [^1] Code is available at https://github.com/berktinaz/stable-concepts.
Published at the 39th Conference on Neural Information Processing Systems (NeurIPS), 2025
@@ 1 Introduction
  • What level of image representation is present in the early stage of the generative process?
@@
Figure 1: General scene layout emerges during the very first generation step in diffusion models. We generate an image with the prompt Several men walking on the dirt with palm trees in the background. Our interpretability framework can predict segmentation masks for each object mentioned in the input prompt, solely relying on model activations cached during the first diffusion step. At this early stage, the posterior mean predicted by the diffusion model does not contain any visual clues about the final generated image.
@@ 2 Background
Interpreting diffusion models – There has been significant effort towards interpreting diffusion models. Tang et al. [49] find that the cross-attention layers in diffusion models with a U-Net backbone – such as SDXL [37] and Stable Diffusion [39] – can be used to generate saliency maps corresponding to textual concepts. Another line of work focuses on finding interpretable editing directions in diffusion U-Nets to control the image generation process. For instance, Kwon et al. [25] and Haas et al. [14] manipulate bottleneck features, Park et al. [35] identifies edit directions based on the SVD of the Jacobian between the input and bottleneck layer of the U-Net, while Chen et al. [5] considers the Jacobian between the input and the posterior mean estimate. Other works modify the key and value projection matrices [34, 11], or directly control object attributes by thresholding attention maps [9, 4].
Beyond editing directions, prior studies have noted the emergence of coarse image structure in the early diffusion steps [27, 15] and the evolution of high- and low-level semantics through different stages of the diffusion process [36, 29]. In contrast, our SAE-based analysis provides a complementary yet more granular and systematic characterization of these phenomena; simultaneously enabling interpretable editing and semantic layout prediction through concept vectors discovered by SAEs.
Recent work [48] trains SAEs on activations of a distilled, single-step diffusion model (SDXL Turbo) [42]. Authors target residual updates in specific cross-attention blocks of the U-Net and found the features in the latent space of SAEs are to be highly interpretable. Our approach differs in two key ways. First, we analyze the time-evolution of interpretable concepts during generation – critical for understanding and controlling diffusion – which single-step models cannot capture. Second, they rely on vision-language foundation models to interpret SAE features by summarizing commonalities among set of images that activate specific features. Alternatively, we introduce a scalable pipeline that builds a flexible concept dictionary that further supplies us with pixel-level annotations using open-set object detectors and segmentation models. Concurrently, Cywiński and Deja [7] apply SAEs to machine unlearning in non-distilled diffusion models. However, they focus on concept removal and train a single SAE across all time steps, whereas we study temporal dynamics by training time-specific SAEs. Lastly, Kim and Ghadiyaram [22] use SAEs for controlled generation, but operate on text-encoder activations rather than visual representations within the diffusion model.
@@ 3.1 SAE Architecture and Loss
In this section, we discuss the design choices behind our SAE model. We opt for $k$-sparse autoencoders (with TopK activation) given their success with GPT-4 [12] and SDXL Turbo [48]. In particular, let $\bm{x}\in\mathbb{R}^{d}$ denote the input activation to the autoencoder that we want to decompose into a sparse combination of features. Then, we obtain the latent $\bm{z}\in\mathbb{R}^{n_{f}}$ by encoding $\bm{x}$ as $\bm{z}=\mathcal{E}\left(\bm{x}\right)=\text{TopK}\left(\text{ReLU}\left(\bm{W}_{enc}\left(\bm{x}-\bm{b}\right)\right)\right)$, where $\bm{W}_{enc}\in\mathbb{R}^{n_{f}\times d}$ denotes the learnable weights of the encoder, $\bm{b}\in\mathbb{R}^{d}$ is a learnable bias term, and TopK function keeps the top $k$ highest activations and sets the remaining ones to 0. Note, that due to the superposition hypothesis, we wish the encoding to be expansive and therefore $n_{f}\gg d$. Then, a decoder is trained to reconstruct the input from the latent $\bm{z}$ in the form $\hat{\bm{x}}=\mathcal{D}\left(\bm{z}\right)=\bm{W}_{dec}\bm{z}+\bm{b}$, where $\bm{W}_{dec}\in\mathbb{R}^{d\times n_{f}}$ represents the learnable weights of the decoder. Note, that the bias term is shared between the encoder and decoder. We refer to $\bm{f}_{i}=\bm{W}_{dec}[:,i]$ columns of $\bm{W}_{dec}$ as concept vectors. We discuss additional details on SAE training in Appendix A.
@@ 3.2 Collecting Model Activations
[…]
To provide an in-depth analysis, we train separate SAEs for different block, conditioning and timestep combinations. Training results are in Appendix A. In this work, we focus on cond features, as we hypothesize that they may be more aligned with human-interpretable concepts due to the direct influence of language guidance through cross-attention (more on this in Appendix C).
@@ 3.3 Extracting Interpretations from SAE Features
In particular, we represent each concept by an associated list of objects, constituting a concept dictionary. The keys are unique concept identifiers (CIDs) assigned to each of the concept vectors of the SAE. The values correspond to objects that commonly occur in areas where the concept is activated. To build the concept dictionary, we first sample a set of text prompts, generate the corresponding images using a diffusion model and extract the SAE activations for each CID during generation. We obtain ground truth annotations for each generated image using a pre-trained vision pipeline, that combines image tagging, object detection and semantic segmentation, resulting in a mask and label for each object in generated images. Finally, we evaluate the alignment between our ground truth masks and the SAE activations for each CID, and assign the corresponding label to the CID only if there is sufficient overlap. We refer the reader to Appendix F.1 for a detailed depiction of the pipeline.
@@ 3.4 Predicting Image Composition from SAE Features
Leveraging the concept dictionary, we predict the final image composition based on SAE features at any time step, allowing us to gain invaluable insight into the evolution of image representations in diffusion models. Suppose that we would like to predict the location of a particular object in the final generated image, but before the reverse diffusion process is completed. First, given SAE features from a given intermediate time step, we extract the top activating concepts for each spatial location. Next, we create a conceptual map of the image by assigning a word embedding to each spatial location based on our curated concept dictionary. This conceptual map shows how image semantics, described by localized word embeddings, vary spatially across the image. Given a concept we would like to localize, such as an object from the input prompt, we produce a target word embedding and compare its similarity to each spatial location in the conceptual map. To produce a predicted segmentation map, we assign the target concept to spatial locations with high similarity, based on a pre-defined threshold value. This technique can be applied to each object present in the input prompt (or to any concepts of interest) to predict the composition of the final generated image. We provide a detailed visualization of our technique in Appendix F.2.
@@ 3.5 Causal Intervention Techniques
Figure 2: An overview of our SAE intervention technique. The prompt "An apple in a basket" specifies the necessary concepts but is vague in terms of spatial composition. We intercept activations of the denoising model and edit the latents after encoding them with the SAE. For the features that are spatially located in the bottom-right quadrant, we increase the coefficient corresponding to "apple" concept, while setting it to 0 for all other features. After the intervention, generated image satisfies the specified layout where all the apples are located in the bottom-right quadrant.
[…]
@@
where $\beta$ is our intervention strength. However, decoding the modified latents directly is suboptimal as the SAE cannot reconstruct the input perfectly. Instead, we modify the activations directly using the concept vectors. The modification in Eq. 1 can be equivalently written as:
@@
An overview of this intervention can be seen in Figure 2. In prior experiments, we observe that the same intervention strength $\beta$ does not work well across different objects $o$. To solve this, we introduce a normalization where the intervention at a spatial coordinate $\left(i,j\right)$ is proportional to the norm of the latent at that coordinate $\left\|\bm{Z}_{\ell,t}[i,j]\right\|$. Therefore, the effective intervention strength is $\beta_{ij}=\beta\left\|\bm{Z}_{\ell,t}[i,j]\right\|$.
Global interventions – Beyond image composition, we investigate whether image style can be manipulated through our discovered concepts. To this end, given a CID $c$ related to the style of interest, as image style is a global property we modify the activation at each spatial location as
@@ 4.1 Building the Concept Dictionary
We sample $40k$ prompts from the LAION-COCO dataset from a split that has not been used to train the SAEs. We build the concept dictionary following our technique introduced in Section 3.3. For annotating generated images, we leverage RAM [53] for image tagging, Grounding DINO [28] for open-set object detection and SAM [23] for segmentation, following the pipeline in Ren et al. [38]. We assign a label to a specific CID if the IoU between the corresponding annotated mask and activation is greater than $0.5$. We binarize the activation map for the IoU calculation by first normalizing to $[0,1]$ range, then thresholding at $0.1$. We visualize the top $5$ activating concepts and the corresponding concept dictionary entries in Appendix F.3.
@@ 4.3 Effectiveness of Interventions Across Diffusion Time
Spatially targeted interventions– We consider bee, book, and dog as the objects of interest and attempt to restrict them to four quadrants: top-left, top-right, bottom-left, and bottom-right. In order to find the CIDs to be intervened on, we sweep the concept dictionary of the given time step and collect all the CIDs where the word of interest appears. Results are summarized in Figure 4.
Global interventions– Through our concept dictionary and visual inspection of top dataset examples at $t=0.5$, we select the following CIDs: #$1722$ that controls the cartoon look of the image, #$524$ appears mostly with beach images where sea and sand are visible together, and #$2137$ activates the most on paintings (top activating images can be found in Appendix H). We find matching concepts for other time steps by picking the CIDs with the highest Word2Vec embedding similarity to the above target CIDs. An overview of results is depicted in Figure 5.